The perimeter of a triangle is 51cm. The length of its sides are consecutive odd integers. Find the length of all three sides.
Answer by macston(5194)

.
S=small side; M=middle side=S+2; L=large side=S+4
.
S+M+F=51cm
S+(S+2)+(S+4)=51cm
3S+6=51cm
3S=45cm
S=15cm
ANSWER 1: The short side is 15cm.
.
M=S+2=15+2=17
ANSWER 2: The middle side is 17cm.
.
L=S+4=15+4=19cm
ANSWER 3: The large side is 19cm.
.